Three years ago Thokozile borrowed R7 500 from Alfred. The condition was that she would pay him back in seven years’ time at an interest rate of 11,21% per year, compounded semi-annually. Six months ago she also borrowed R25 000 from Alfred at 9,45% per year, compounded monthly. Thokozile would like to pay off her debt four years from now


The amount of money that Thokozile will have to pay Alfred four years from now is

[1] R36 607,98.

[2] R45 181,81.

[3] R55 336,49.

[4] R48 032,20.

[5] R54 278,92


1
Expert's answer
2021-07-26T09:43:36-0400

"P=7500"

A=7500+7500(1+0.1121/2)14

A=23593.25

P'=25000

A'=25000+25000(1+0.0945/12)54

Total deposited=deposited=23593.25+63185.66=86778.91